OpenStudy (anonymous):

convert inches to meter 120inches squared to m squared equals what

OpenStudy (anonymous):

can anyone help me solve this

OpenStudy (anonymous):

1 inch = 2.54 cms 1 inch squared = 2.54^2 sq cms 120 sq ins = 120 * 2.54^2 sq cms

OpenStudy (anonymous):

divide this by 10000 to get sq meters